OpenStudy (anonymous):

Write the equation of a parabola with a vertex at (10, 0) and a directrix x = 8.

OpenStudy (anonymous):

x+10= 1/8^x2 y-10= 1/8^x2 y+10= 1/8^x2 x-10= 1/8^x2 are my choices

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

anything? @baad1994

OpenStudy (anonymous):

 The distance from V to the directrix is the same from V to F So you have this: P=-2 (Negative because the parabola is open to the left) and V(8,0) The form of your equation is \[(y-k)^{2}=4p(x-h)\] where (h,k) is your V and p is the distance from V to F \[(y-0)^{2}=-4*2(x-8)\] \[(y-0)^{2}=-8(x-8)\] \[y^{2}=-8x+64\]

OpenStudy (anonymous):

the image is wrong .. is this image